### v1

# VocabTailor: Dynamic Vocabulary Selection for Downstream Tasks in Small Language Models

Abstract
Small Language Models (SLMs) provide computational advantages in resource-constrained environments, yet memory limitations remain a critical bottleneck for edge device deployment. A substantial portion of SLMs’ memory footprint stems from vocabulary-related components, particularly embeddings and language modeling (LM) heads, due to large vocabulary sizes. Existing static vocabulary pruning, while reducing memory usage, suffers from rigid, one-size-fits-all designs that cause information loss from the prefill stage and a lack of flexibility. In this work, we identify two key principles underlying the vocabulary reduction challenge: the lexical locality principle, the observation that only a small subset of tokens is required during any single inference, and the asymmetry in computational characteristics between vocabulary-related components of SLM. Based on these insights, we introduce VocabTailor, a novel decoupled dynamic vocabulary selection framework that addresses memory constraints through offloading embedding and implements a hybrid static-dynamic vocabulary selection strategy for LM Head, enabling on-demand loading of vocabulary components. Comprehensive experiments across diverse downstream tasks demonstrate that VocabTailor achieves a reduction of up to 99% in the memory usage of vocabulary-related components with minimal or no degradation in task performance, substantially outperforming existing static vocabulary pruning.
1 Introduction
Large Language Models (LLMs) (Brown et al. 2020; Touvron et al. 2023a, b; Achiam et al. 2023; Team et al. 2024; Anthropic 2024, 2025; Bai et al. 2023; Guo et al. 2025) have rapidly become foundational to modern AI applications. Recently, increasing attention has turned towards small language models (SLMs), which are better suited for deployment on edge devices and in resource-constrained environments. Despite their compact size, memory still remains a bottleneck, particularly for edge devices with limited GPU memory. A key driver of this bottleneck is the model’s vocabulary size, which directly impacts the memory footprint of both the embedding layer and the language modeling (LM) head. For example, in the Llama 3.2 1B model with a 128K-token vocabulary, the embedding and LM head account for over 20% of the total memory usage. As SLMs are scaled down and deployed under tighter memory constraints, vocabulary-related memory inefficiencies become increasingly unsustainable, posing a fundamental barrier to efficient SLM deployment.
To address this, prior work has explored static vocabulary pruning strategies (Ushio, Zhou, and Camacho-Collados 2023; Yang, Cui, and Chen 2022), which reduce vocabulary size by eliminating rare or irrelevant tokens based on curated corpora. While these approaches are well-motivated, they suffer from key limitations due to their static and coupled design, which assumes a single, globally pruned vocabulary is applied on all vocabulary-related components (i.e.,  tokenizer, embedding, and LM head). This design introduces two major issues:
Figure 1: Overview of VocabTailor
  • Premature Information Loss: Pruning the tokenizer, embedding, and LM head altogether alters the input representation passed to the transformer. These modified inputs may differ significantly from those seen during pretraining, causing distributional shifts and information loss from the prefill stage. Notably, each pruning step in the pipeline introduces information loss that accumulates, leading to cumulative performance degradation during inference.
  • Lack of flexibility and Adaptability: The static strategy limits the model’s adaptability across diverse tasks. Supporting different task configurations typically requires duplicating multiple copies of vocabulary-related components, resulting in substantial storage overhead and increased complexity in deployment.
Based on empirical observations and theoretical analysis, we derive two key principles for efficient design:
  • Lexical Locality: In common downstream tasks, generation relies on a highly localized vocabulary—each output depends on a small subset of input tokens and a limited set of task-specific tokens.
  • Computation Asymmetry: There is computation asymmetry between embedding and LM head. The embedding layer mainly leverages the lookup operation, which is computationally cheap but limited by memory bandwidth. In contrast, the LM head is compute intensive with massive matrix multiplication requiring immense floating point power, making it an ideal task for a GPU. Existing pruning methods ignore such asymmetry, missing opportunities for system-level optimization.
Guided by these principles, we propose VocabTailor, a flexible and efficient framework for dynamic vocabulary selection. VocabTailor is based on two main pillars:
  • Decoupled Design: We adopt a decoupled design for vocabulary-related components. We retain the full tokenizer and offload the embedding layer to CPU memory. Since the embedding lookup is a memory-intensive operation with $\mathcal{O(1)}$ computational complexity, we can strategically offload it to free up valuable GPU memory, with minimal overhead to the overall system performance.
  • Hybrid Static-Dynamic Vocabulary Selection: At runtime, we dynamically select and load input-relevant tokens while maintaining a small, static set of task-specific tokens to ensure stable and efficient computation in the LM head.
This design enables substantial memory savings without compromising input fidelity and model generality. Compared to static pruning, which retains the union of all input and output tokens $\left(\bigcup\mathcal{I}_{i}\right)\bigcup\left(\bigcup\mathcal{O}_{i}\right)$, VocabTailor only needs $\mathcal{I}_{i}\bigcup\mathcal{T}$ at inference time, where $\mathcal{I}_{i}$, $\mathcal{O}_{i}$ are input and output tokens for example $i$, and $\mathcal{T}$ is a small, fixed task-specific token set. Since $\bigcup\mathcal{I}_{i}\gg\mathcal{I}_{i}$, this leads to substantial memory savings and improved task adaptability.
In summary, this paper makes the following contributions:
  1. 1.
    We present the first systematic analysis of vocabulary management in LLMs through the lens of lexical locality and computation asymmetry.
  2. 2.
    We propose VocabTailor, a flexible, memory-efficient, and task-adaptive framework that supports a hybrid static-dynamic vocabulary selection strategy along with an enhanced profiling strategy.
  3. 3.
    Across five representative downstream tasks—machine translation, summarization, code completion, information extraction, and math problem solving—VocabTailor reduces memory usage of vocabulary-related components by up to 99%, significantly lowers memory usage with minimal or no performance degradation.
2 Related Work
2.1 Small Language Model
Small language models (SLMs) are compact alternatives to large language models (LLMs), which are designed for efficiency, lower computational costs, and deployment on resource-constrained devices. While LLMs like GPT-4 (Achiam et al. 2023), LLaMA (Touvron et al. 2023a, b), Claude (Anthropic 2024, 2025), Gemini (Team et al. 2024), Qwen (Bai et al. 2023), and DeepSeek (Guo et al. 2025) have achieved widespread success across real-life applications, SLMs are gaining attention due to their suitability in GPU memory-constrained environments, personal devices, and task-specific scenarios. They offer a scalable, efficient, and sustainable solution tailored for real-time and on-device applications (Lamaakal et al. 2025). With careful selection, small open models can rival and even outperform LLMs while offering improved speed and memory efficiency (Sinha, Jain, and Chadha 2024).
Despite advances in architectures, training techniques, and model compression techniques, SLMs still face challenges, including trade-offs between model size and accuracy, generalization limitations, and concerns over bias and privacy (Van Nguyen et al. 2024; Lamaakal et al. 2025). A common approach to building SLMs is distilling them from LLMs while retaining the same tokenizer and vocabulary. This typically causes vocabulary-related components (i.e.,  embedding and LM Head) to account for a large proportion of the model’s total parameters. This makes vocabulary pruning an effective optimization strategy for SLMs.
2.2 Tokenization
Tokenization is a fundamental preprocessing step in Natural Language Processing (NLP) that splits text into smaller units called tokens (e.g., words or characters), which form the input to downstream tasks. Over the years, various tokenization techniques have been developed (Sennrich, Haddow, and Birch 2016; Kudo and Richardson 2018; Devlin et al. 2019). Among these, Byte-Pair Encoding (BPE) (Sennrich, Haddow, and Birch 2016) has become one of the most widely used. Originally developed for data compression, BPE was adapted to tokenize text by iteratively merging the most frequent adjacent symbol pairs starting from individual characters until a target vocabulary size is reached. This approach enables efficient representation of frequent words with fewer tokens while breaking rare or unseen words into informative subword units. BPE’s widespread adoption across transformer architectures has established it as a core component of LLM infrastructure.
State-of-the-art LLMs such as GPT-4 (Achiam et al. 2023), LLaMA (Touvron et al. 2023a, b), Gemini (Team et al. 2024), Claude (Anthropic 2024, 2025), and DeepSeek (Guo et al. 2025) rely on BPE-based tokenizers to balance vocabulary efficiency and expressiveness. However, the resulting vocabulary sizes are often large, leading to large embedding matrices and LM heads, which increase computational and memory overhead during inference. This scalability bottleneck has motivated research into vocabulary reduction and adaptive tokenization strategies, especially for resource-constrained deployments.
Importantly, pruning a BPE-based tokenizer can interfere with its learned merge operations governing subword segmentation, resulting suboptimal tokenizations. This disruption reduces the granularity of lexical representations, ultimately impairing model performance on downstream tasks.
2.3 Vocabulary Pruning
Vocabulary pruning has emerged as a key area of research in NLP, particularly for scaling and deploying efficient language models. During the BERT era, interest in pruning surged as researchers explored ways to streamline BERT and other transformer-based models. More recently, with the rise of small language models (SLMs), efficient vocabulary selection has once again become a pressing concern.
Vocabulary-trimming (VT) (Ushio, Zhou, and Camacho-Collados 2023) and TextPruner (Yang, Cui, and Chen 2022) were proposed to improve efficiency and reduce model size by removing tokens irrelevant to the target language or rarely seen in downstream tasks. Both methods follow a static pruning strategy: they identify language-specific or task-relevant tokens from a curated corpus and prune the vocabulary accordingly. While this process simultaneously reduces the size of the tokenizer, embedding, and LM head, it introduces premature information loss, leading to cumulative performance degradation, and limited flexibility and adaptability, resulting in substantial memory overhead and increased deployment complexity. These two major limitations underscore the need for more flexible and dynamic vocabulary pruning strategies that are task-specific, adaptable, and generalizable across deployment settings. In response, we propose our framework in Section 3.
3 Method
3.1 VocabTailor Framework
As illustrated in Fig. 1, VocabTailor decouples vocabulary management across vocabulary-related components. Unlike static vocabulary pruning that uniformly reduces the tokenizer, embedding layer, and LM head, VocabTailor treats each component based on its unique computational and storage properties.
For the tokenizer, we retain the full vocabulary to preserve input fidelity and prevent information loss during the prefill stage. This avoids the cumulative performance degradation caused by pruned tokenizers, which often fail to fully capture the input expressiveness.
For the embedding layer and LM head, despite their similar roles(tokenize and detokenize), a fundamental computation asymmetry exists between them.
For the embedding layer, which relies on simple lookup operations, is naturally CPU-friendly and incurs negligible runtime overhead when offloaded to CPU memory from high-cost accelerator memory (Yu et al. 2025). Thus, we retain it fully and offload it to CPU.
In contrast, for the LM head, which performs compute-intensive matrix multiplications, must remain on GPU for efficient inference. To optimize its memory footprint, we introduce a hybrid static-dynamic vocabulary selection strategy. This approach leverages the lexical locality of tokens and significantly reduces memory usage while preserving downstream task performance and enabling flexible, task-specific adaptation without model duplication.
This decoupled architecture addresses the key limitations of existing approaches while providing a theoretically grounded and practically efficient solution for large-scale vocabulary management. The detailed design and implementation of vocabulary selection strategy are presented in Sections 3.23.4.
Figure 2: Left: Input-output lexical overlap ratio. Right: Example of input-output lexical overlap in summarization task.
3.2 Motivation: Analysis of Lexical Locality
To investigate the essence of efficient vocabulary selection, we analyze input-output pairs across diverse datasets aligned with various downstream tasks. Our empirical analysis reveals two fundamental properties of lexical locality that motivate VocabTailor’s dynamic vocabulary selection strategy.
Observation 1
Input-Driven Locality: Common downstream tasks exhibit strong input-output lexical overlap—each output contains a small subset of input tokens.
In various downstream tasks, the model output reuses tokens from the input, as shown in Fig. 2. This phenomenon is particularly pronounced in text extraction tasks (e.g., span-based QA or named entity recognition), where output tokens are typically a subset of the input tokens. In code completion, generated code often replicates variable names, function names, and other identifiers from the input context. Similarly, in text summarization, the generated summary contains large spans from the source document. For instance, in summarization, on average 61.9% of tokens in generated summaries are copied from the input document. From an information-theoretic perspective, the input context dramatically reduces the entropy of the output token distribution, constraining generation to a much smaller effective vocabulary. Thus, preserving input vocabulary is critical for maintaining performance during vocabulary reduction.
Observation 2
Task-Driven Locality: Vocabulary required for generation is highly localized—each output depends on a limited set of task-specific tokens.
Due to input diversity across datasets, the union of all input tokens is significantly larger than the tokens required for any single generation instance. Each input introduces unique tokens, making the aggregate input vocabulary much larger than individual requirements. Crucially, the remaining output tokens—those not found in the corresponding input—form a relatively small, task-specific set $\mathcal{T}$ that captures the essential generation patterns for the task.
Formally, let $\mathcal{I}$ denote the set of all input tokens, $\mathcal{I}_{i}$ the set of input tokens for instance $i$ in a dataset, $\mathcal{O}$ the set of all output tokens, and $\mathcal{T}\subset\mathcal{O}$ the set of all task-specific tokens in outputs. We observe that:
$\mathcal{I}=\bigcup_{i}\mathcal{I}_{i}\gg\mathcal{I}_{i}\;\text{and }\mathcal{O}>\mathcal{T}$
These observations expose the fundamental inefficiency of static vocabulary pruning that operates on the union $\mathcal{I}\cup\mathcal{O}$ for every example. VocabTailor exploits this lexical locality by dynamically constructing active per-example vocabularies using only per-example input tokens $\mathcal{I}_{i}$ and a compact task-specific set $\mathcal{T}$:
$\mathcal{I}\bigcup\mathcal{O}\gg\mathcal{I}_{i}\bigcup\mathcal{T}$
This dynamic targeting of the much smaller $\mathcal{I}_{i}\cup\mathcal{T}$ enables substantial memory savings without compromising generation quality, as it only retains the tokens necessary for each inference instance.
3.3 The Hybrid Vocabulary Selection Strategy
VocabTailor utilizes lexical locality through a hybrid architecture that combines dynamic runtime adaptation with static offline optimization, enabling efficient vocabulary management without sacrificing generation flexibility.
Dynamic Selection (Runtime Behavior)
At the start of each inference instance, VocabTailor identifies the unique input token set $\mathcal{I}_{i}$ from the input text and selectively loads the corresponding LM head weight vectors from CPU main memory to GPU. This selective loading exploits the input-driven locality principle, ensuring that only input-relevant vocabulary components are active during generation. Crucially, we overlap this memory transfer with the model’s forward propagation, leveraging GPU-CPU parallelism to mitigate runtime overhead.
Static Selection (Offline Construction)
The static component maintains a compact, task-specific core vocabulary $\mathcal{T}$ that captures essential output tokens independent of input context. While simple frequency-based filtering proves inadequate due to noise from typos and multilingual interference in broad corpora, we introduce a theoretically grounded filtering pipeline (Algorithm 1) that constructs $\mathcal{T}$ through fine-grained analysis.
Algorithm 1 Profiling Strategy
Definition:
$\mathcal{V}$: Corpus-profiling vocabulary
$M$: Total number of documents in the corpus
$\mathcal{I}_{i}$: Set of input tokens corresponding to a single example $i$
$\mathcal{U}$: Set of language-specific tokens obtained from Unicode blocks
$df\in\mathbb{R}$: Dictionary mapping token $v$ to its document frequency $df(v)$
$\tau\in[0,1]$: Tolerance threshold (fraction of documents allowed to be impacted)
$\mathcal{T}$: Final calibrated task-specific vocabulary
Input: $\mathcal{V}$, $\tau$
Output: $\mathcal{T}$
1:Input-Aware Filtering
[…]
3:Language-specific Filtering
[…]
5:Tolerance Filtering
[…]
8: Sort $F$ ascending by $df(v)$
[…]
10:for $j\leftarrow 1$ to $N+1$ do
11:  if sum($F[:\texttt{j}])>\tau M$ then
12:   $\texttt{index}\leftarrow N-j+1$
13:   break
14:  end if
15:end for
16:
17:return  $\mathcal{T}\leftarrow\{v\in F[-\texttt{index}:]\}$
3.4 Fine-grained Construction of the Static Task Vocabulary
Our static vocabulary construction algorithm addresses the challenge of isolating truly task-essential tokens through a three-stage filtering process:
Input-Aware Filtering
We first exclude all input tokens from the candidate vocabulary, isolating tokens the model must generate without input cues (e.g., function keywords in code generation, discourse markers in summarization). This step directly implements task-driven locality by identifying the irreducible core $\mathcal{T}$ that cannot be derived from the input context.
Language-Specific Filtering
To suppress noise in mixed-language scenarios (e.g., code datasets with multilingual comments), we apply Unicode block analysis to retain only tokens belonging to the target language family. This heuristic-based approach effectively handles cross-lingual interference while preserving task-relevant vocabulary.
Tolerance Filtering
When additional vocabulary reduction is required, we formulate vocabulary selection similar to a pruning problem. We define the impact metric $df(v)$ as the document frequency—the fraction of instances where token $v$ appears in the ground truth. Tokens are sorted by ascending $df(v)$ and iteratively removed until the cumulative document frequency of pruned tokens reaches a user-defined tolerance threshold $\tau$. This threshold bounds the worst-case performance drop, where $\tau=0.01$ ensures that at most 1% of profiling samples lose a critical token.
VocabTailor’s hybrid approach bridges theoretical rigor (exploiting lexical locality principles) with practical efficiency (optimizing GPU-CPU memory hierarchy). The static-dynamic decomposition enables substantial memory savings while preserving the model’s generational capabilities, providing a scalable solution for deploying SLMs across diverse applications with controllable efficiency-accuracy trade-offs.
4 Experiments
4.1 Settings
Tasks and datasets
[…]
Evaluation metrics
We include sacreBLEU (Post 2018), METEOR (Banerjee and Lavie 2005), and COMET (Rei et al. 2020) for machine translation. For summarization, we use Rouge-1, Rouge-2, and Rouge-L scores (Lin 2004). Pass@1 (Chen et al. 2021) is used for code completion. We use F1 score for information extraction, and accuracy for math problem solving. These metrics are standard in the field and provide a robust measure of model performance across the target task.
Models
For machine translation, we use Qwen3-1.7B (Yang et al. 2025). For summarization and information extraction, we employ Llama 3.2 1B (Dubey et al. 2024). For summarization, the base model is fine-tuned for a better base performance. For code completion, we choose deepseek-coder-1.3b-base (Guo et al. 2024), and for math problem solving, we apply rho-math-1b-interpreter-v0.1 (Lin et al. 2024).
Baselines and Other Settings
We compare our method (VocabTailor) with the original model (Original) and static vocabulary pruning (VP). Static vocabulary pruning follows the common routines of corpus-based filtering. For a fair comparison, VP and VocabTailor use the same profiling corpra. We set the tolerance threshold $\tau=0.01$ on all tasks for VocabTailor. As a hybrid dynamic-static framework, the vocabulary size of VocabTailor model varies in each single inference. Here we report the average vocabulary size for each downstream task. For all models, we set the temperature to 0 to avoid randomness.
4.2 Results
Task Model Vocabulary Metric
[…]
VP 64,117 (42.28%) 14.840/11.842/81.191
VocabTailor (Ours) 18,874 + [40] (12.47%) 15.393/12.692/81.440
[…]
VP 65,518 (43.21%) 22.585/46.735/74.933
VocabTailor (Ours) 24,185 + [14] (15.96%) 21.127/46.970/75.493
Summarization Original 128,000 0.364/0.147/0.292
VP 59,613 (90.99%) 0.364/0.147/0.292
VocabTailor (Ours) 36,332 + [15] (55.48%) 0.364/0.147/0.292
Code Completion Original 32,000 54.097%
VP 24,888 (77.78%) 8.124%
VocabTailor (Ours) 3,521 + [58] (11.18%) 53.865%
Information Extraction Original 128,000 38.400
VP 49,106 (38.36%) 2.580
VocabTailor (Ours) [105] (0.08%) 62.730
Math Problem Solving Original 32,000 88.40
VP 10,300 (32.19%) 87.80
VocabTailor (Ours) 5,135 + [14] (16.09%) 88.40
Table 1: Results on machine translation (sacreBLEU/METEOR/COMET), summarization (Rouge-1/Rouge-2/Rouge-L), code completion (Pass@1), information extraction (F-1), and math problem solving (Accuracy), including the vocabulary size and the ratio to the original model (%). For VocabTailor, vocabulary size consists of task-specific tokens and the average number of dynamic tokens highlighted in brackets. The best results are in bold characters.
As shown in Table 1, VocabTailor consistently achieves substantial vocabulary reduction while maintaining or even improving task performance compared with the original model, outperforming static pruning in nearly all cases.
Machine Translation
[…]
Summarization
For summarization, all three models yield identical ROUGE scores, indicating no difference in summary quality. However, VocabTailor achieves this with just 55% of the original vocabulary, 35% fewer tokens than VP. These results demonstrate that aggressive vocabulary reduction is possible without compromising output quality when apply a hybrid dynamic-static vocabulary selection.
Code Completion
On the SAFIM benchmark with deepseek-coder-1.3b-base, VP leads to a dramatic drop in performance: Pass@1 falls from 54.097% to 8.124%, despite a relatively modest vocabulary reduction (down to 77.78% of the original size). This highlights a key limitation of corpus-based pruning for code: essential elements such as variable names and identifiers may be discarded if they appear infrequently in the profiling corpus. In contrast, VocabTailor retains only 11.18% of the vocabulary and yet achieves a high Pass@1 of 53.865%. These findings underscore the robustness of our input- and task-specific vocabulary selection strategy for generation-intensive tasks such as code synthesis.
Information Extraction
For information extraction tasks, we evaluate on the SQuAD dataset using LLaMA 3.2 1B. VocabTailor achieves the most striking result: with just 0.08% of the original vocabulary retained (and no static tokens at all), it attains an F1 score of 62.73, outperforming both the Original (38.4) and VP (2.58) by large margins. This result stems from the nature of extractive tasks, where the output vocabulary is typically a subset of the input. Because VocabTailor preserves input tokens dynamically, it retains all the necessary vocabulary for accurate extraction. The poor performance of VP suggests that static, corpus-based pruning is poorly suited for tasks where input-output overlap is high, whereas VocabTailor is especially effective.
Math Word Problem Solving
For symbolic reasoning, we evaluate on a math problem-solving task. The unpruned model achieves a score of 88.40, which is fully preserved under VocabTailor, even though it reduces the vocabulary to just 16% of the original. VP, while also reducing vocabulary size to 32%, causes a slight performance drop to 87.80. This shows that even in tasks requiring high-precision symbolic handling, our approach remains effective.
Across five distinct tasks, VocabTailor consistently demonstrates strong performance while substantially reducing vocabulary size. In many cases, it matches or even exceeds the performance of the unpruned model. Compared to static pruning, which often compromises accuracy, VocabTailor reduces vocabulary more aggressively (up to 99%) without sacrificing model quality. These findings support that the dynamic vocabulary selection is a practical and efficient approach to deploying SLMs in resource-constrained environments.
4.3 Ablation Study
To understand the contributions of each component in our framework, we conduct a series of ablation experiments on the SAFIM dataset using the deepseek-coder-1.3b-base model for the code completion task. We primarily focus on evaluating the impact of different vocabulary configurations, including the static and dynamic components, as well as our proposed three-stage filtering process.
Impact of Dynamic vs. Static Vocabulary Components
We first evaluate the impact of the dynamic and static vocabulary components, both individually and in combination. As shown in Table 2, using only the dynamic part that contains tokens profiled from the specific input examples, results in a significant performance drop (Pass@1 of 36.064%). This demonstrates that input tokens alone lack the broader coverage needed for robust code completion. Using only the static part (task-specific tokens) achieves a Pass@1 of 52.3%. However, this still underperforms the full static-dynamic configuration, which achieves the best result at 53.865% with nearly the same vocabulary size. This indicates that while the static tokens carry most of the task-relevant capacity, including the dynamic tokens adds critical input-specific nuances, and their combination is essential for optimal performance.
We also compare our static-only approach with VP: VP retains 78% of the original vocabulary—more than our approach—it results in a drastically lower Pass@1 of 8.124%. This stark contrast underscores a key insight: while both VP and our static-only setup are static, VP’s direct modification of the tokenizer and embeddings damages input representations, leading to severe degradation. In contrast, our method retains the full tokenizer and embedding, thereby preserving representational integrity and maintaining high performance even with significantly fewer tokens.
Model Vocabulary Pass@1
Original 100% 54.097%
VP 77.80% 8.124%
Dynamic + Static ($\tau=0.01$) 11.18% 53.865%
Dynamic only 0.81% 36.064%
Static only ($\tau=0.01$) 11.00% 52.300%
Table 2: Comparison of the dynamic and static components in VocabTailor.
Impact of Input-aware and Language-specific Filtering
As discussed earlier, VocabTailor calibrates the static vocabulary through three-stage filtering. As shown in Table 3, starting from the unfiltered static vocabulary (78% of the original), applying input-aware filtering (IA) reduces the size to 53% with virtually no performance loss (54.074% vs. 54.062%). Adding language-specific filtering (LS) further reduces the vocabulary size to 46%, while performance slightly improves to 54.085%. This improvement likely stems from the removal of noisy or irrelevant tokens from multilingual corpora, allowing the model to focus on task-relevant representations. These results demonstrate that IA and LS can significantly compress the vocabulary without degrading accuracy, validating the effectiveness of our static token selection process.
Model Vocabulary Pass@1
Original 100.00% 54.097%
Dynamic + Unfiltered static 77.78% 54.074%
Dynamic + IA 52.85% 54.062%
Dynamic + IA + LS 45.61% 54.085%
Table 3: Comparison of input-aware filtering (IA) and language-specific filtering (LS) in VocabTailor.
Impact of Tolerance Filtering
Lastly, we analyze the effect of tolerance filtering, which allows for further reduction of rarely activated tokens based on cumulative document frequency. In Table 4, we vary the tolerance threshold ($\tau$) to observe the trade-off between vocabulary size and accuracy. At $\tau=0$, we preserve all profiled tokens, achieving a Pass@1 of 54.085%. As tolerance increases, more tokens are filtered out, reducing the vocabulary to as low as 2.5% of the original size ($\tau=0.10$), with a gradual decline in performance. Importantly, even with $\tau=0.01$, the vocabulary shrinks to 11.18% with only 1.8% drop in Pass@1, suggesting that our method is robust to aggressive reduction. This highlights the flexibility of tolerance as a tuning knob to balance compression vs. accuracy.
Model Vocabulary Pass@1
$\tau=0$ 45.61% 54.085%
$\tau=0.01$ 11.18% 53.865%
$\tau=0.02$ 7.28% 53.714%
$\tau=0.10$ 2.50% 52.277%
Table 4: Comparison of different tolerance thresholds in VocabTailor.
5 Conclusion
In this paper, we propose a flexible and efficient vocabulary selection framework effectively reduce memory usage during SLM inference. By identifying and leveraging lexical locality together with the computation asymmetry, our method reduce up to 99% in the memory usage of vocabulary-related components of SLM while maintain the performance on representative downstream tasks.
6 Limitations & Future Work
The proposed framework presented in this paper is only explored on language downstream tasks of SLMs. This method may be extended and applied to VLMs, ALMs, and MLLMs for memory-efficient inference on image/video understanding and audio generation tasks. Custom kernel implementations and hardware-specific optimizations can be integrated into the framework for further inference optimization.

### v2

@@
 Comprehensive experiments across diverse downstream tasks demonstrate that VocabTailor achieves a reduction of up to 99% in the memory usage of vocabulary-related components with minimal or no degradation in task performance, substantially outperforming existing static vocabulary pruning.
+VocabTailor: Dynamic Vocabulary Selection for Downstream Tasks in Small Language Models
+Hanling Zhang1,2thanks: Equal contribution., Yayu Zhou611footnotemark: 1, Tongcheng Fang 3, Zhihang Yuan3thanks: Co-corresponding author., Guohao Dai 1,5, Wanli Ouyang 2,4, Yu Wang 322footnotemark: 2 1Infinigence AI, 2The Chinese University of Hong Kong, 3Tsinghua University, 4Shanghai Artificial Intelligence Laboratory, 5Shanghai Jiao Tong University, 6Independent Researcher
 1 Introduction
-Large Language Models (LLMs) (Brown et al.
-2020; Touvron et al.
-2023a, b; Achiam et al.
-2023; Team et al.
-2024; Anthropic 2024, 2025; Bai et al.
-2023; Guo et al.
-2025) have rapidly become foundational to modern AI applications.
+Figure 1: Overview of VocabTailor
+Large Language Models (LLMs) (brown2020language; touvron2023llama; touvron2023llama-1; achiam2023gpt; team2024gemini; anthropic_claude_2024; anthropic_introducing_2025; bai2023qwen; guo2025deepseek) have rapidly become foundational to modern AI applications.
 Recently, increasing attention has turned towards small language models (SLMs), which are better suited for deployment on edge devices and in resource-constrained environments.
@@
 As SLMs are scaled down and deployed under tighter memory constraints, vocabulary-related memory inefficiencies become increasingly unsustainable, posing a fundamental barrier to efficient SLM deployment.
-To address this, prior work has explored static vocabulary pruning strategies (Ushio, Zhou, and Camacho-Collados 2023; Yang, Cui, and Chen 2022), which reduce vocabulary size by eliminating rare or irrelevant tokens based on curated corpora.
+To address this, prior work has explored static vocabulary pruning strategies (ushio_efficient_2023; yang_textpruner_2022), which reduce vocabulary size by eliminating rare or irrelevant tokens based on curated corpora.
 While these approaches are well-motivated, they suffer from key limitations due to their static and coupled design, which assumes a single, globally pruned vocabulary is applied on all vocabulary-related components (i.e.,  tokenizer, embedding, and LM head).
-This design introduces two major issues:
-Figure 1: Overview of VocabTailor
-  • Premature Information Loss: Pruning the tokenizer, embedding, and LM head altogether alters the input representation passed to the transformer.
+This design introduces two major issues.
+First, premature information loss arises because pruning the tokenizer, embedding, and LM head altogether alters the input representation passed to the transformer.
 These modified inputs may differ significantly from those seen during pretraining, causing distributional shifts and information loss from the prefill stage.
 Notably, each pruning step in the pipeline introduces information loss that accumulates, leading to cumulative performance degradation during inference.
-  • Lack of flexibility and Adaptability: The static strategy limits the model’s adaptability across diverse tasks.
+Second, lack of flexibility and adaptability as the static strategy limits the model’s adaptability across diverse tasks.
 Supporting different task configurations typically requires duplicating multiple copies of vocabulary-related components, resulting in substantial storage overhead and increased complexity in deployment.
-Based on empirical observations and theoretical analysis, we derive two key principles for efficient design:
-  • Lexical Locality: In common downstream tasks, generation relies on a highly localized vocabulary—each output depends on a small subset of input tokens and a limited set of task-specific tokens.
-  • Computation Asymmetry: There is computation asymmetry between embedding and LM head.
-The embedding layer mainly leverages the lookup operation, which is computationally cheap but limited by memory bandwidth.
-In contrast, the LM head is compute intensive with massive matrix multiplication requiring immense floating point power, making it an ideal task for a GPU.
+Based on empirical observations and theoretical analysis, we derive two key principles for efficient design.
+Lexical locality captures the empirical observations that, in common downstream tasks, generation relies on a highly localized vocabulary, where each output depends on a small subset of input tokens and a limited set of task-specific tokens.
+Computation asymmetry reflects the distinct computational characteristics of the embedding and LM head: the embedding layer mainly leverages lookup operations, which are computationally cheap but memory-bandwidth bound, whereas the LM head is compute-intensive with massive matrix multiplication requiring immense floating-point power and thus better suited for GPUs.
 Existing pruning methods ignore such asymmetry, missing opportunities for system-level optimization.
-Guided by these principles, we propose VocabTailor, a flexible and efficient framework for dynamic vocabulary selection. VocabTailor is based on two main pillars:
-  • Decoupled Design: We adopt a decoupled design for vocabulary-related components.
+Guided by these principles, we propose VocabTailor, a flexible and efficient framework for dynamic vocabulary selection.
+VocabTailor is based on two main pillars.
+First, we adopt a decoupled design for vocabulary-related components.
 We retain the full tokenizer and offload the embedding layer to CPU memory.
 Since the embedding lookup is a memory-intensive operation with $\mathcal{O}(1)$ computational complexity, we can strategically offload it to free up valuable GPU memory, with minimal overhead to the overall system performance.
-  • Hybrid Static-Dynamic Vocabulary Selection: At runtime, we dynamically select and load input-relevant tokens while maintaining a small, static set of task-specific tokens to ensure stable and efficient computation in the LM head.
+Second, we propose hybrid static-dynamic vocabulary selection: at runtime, we dynamically select and load input-relevant tokens while maintaining a small, static set of task-specific tokens to ensure stable and efficient computation in the LM head.
 This design enables substantial memory savings without compromising input fidelity and model generality.
 Compared to static pruning, which retains the union of all input and output tokens $\left(\bigcup\mathcal{I}_{i}\right)\bigcup\left(\bigcup\mathcal{O}_{i}\right)$, VocabTailor only needs $\mathcal{I}_{i}\bigcup\mathcal{T}$ at inference time, where $\mathcal{I}_{i}$, $\mathcal{O}_{i}$ are input and output tokens for example $i$, and $\mathcal{T}$ is a small, fixed task-specific token set.
-Since $\bigcup\mathcal{I}_{i}\gg\mathcal{I}_{i}$, this leads to substantial memory savings and improved task adaptability.
-In summary, this paper makes the following contributions:
+Since $|\bigcup\mathcal{I}_{i}|\gg|\mathcal{I}_{i}|$, this leads to substantial memory savings and improved task adaptability.
+In summary, our contributions are:
   1. 1.
@@
   3. 3.
-    Across five representative downstream tasks—machine translation, summarization, code completion, information extraction, and math problem solving—VocabTailor reduces memory usage of vocabulary-related components by up to 99%, significantly lowers memory usage with minimal or no performance degradation.
+    Across five representative downstream tasks—machine translation, summarization, code completion, information extraction, and math problem solving—VocabTailor reduces memory usage of vocabulary-related components by up to 99%, with minimal or no performance degradation.
 2 Related Work
@@
 Small language models (SLMs) are compact alternatives to large language models (LLMs), which are designed for efficiency, lower computational costs, and deployment on resource-constrained devices.
-While LLMs like GPT-4 (Achiam et al.
-2023), LLaMA (Touvron et al.
-2023a, b), Claude (Anthropic 2024, 2025), Gemini (Team et al.
-2024), Qwen (Bai et al.
-2023), and DeepSeek (Guo et al.
-2025) have achieved widespread success across real-life applications, SLMs are gaining attention due to their suitability in GPU memory-constrained environments, personal devices, and task-specific scenarios.
-They offer a scalable, efficient, and sustainable solution tailored for real-time and on-device applications (Lamaakal et al.
-2025).
-With careful selection, small open models can rival and even outperform LLMs while offering improved speed and memory efficiency (Sinha, Jain, and Chadha 2024).
-Despite advances in architectures, training techniques, and model compression techniques, SLMs still face challenges, including trade-offs between model size and accuracy, generalization limitations, and concerns over bias and privacy (Van Nguyen et al.
-2024; Lamaakal et al.
-2025).
+While LLMs like GPT-4 (achiam2023gpt), LLaMA (touvron2023llama; touvron2023llama-1), Claude (anthropic_claude_2024; anthropic_introducing_2025), Gemini (team2024gemini), Qwen (bai2023qwen), and DeepSeek (guo2025deepseek) have achieved widespread success across real-life applications, SLMs are gaining attention due to their suitability in GPU memory-constrained environments, personal devices, and task-specific scenarios.
+They offer a scalable, efficient, and sustainable solution tailored for real-time and on-device applications (lamaakal2025tiny).
+With careful selection, small open models can rival and even outperform LLMs while offering improved speed and memory efficiency (sinha2024small).
+Despite advances in architectures, training techniques, and model compression techniques, SLMs still face challenges, including trade-offs between model size and accuracy, generalization limitations, and concerns over bias and privacy (van2024survey; lamaakal2025tiny).
 A common approach to building SLMs is distilling them from LLMs while retaining the same tokenizer and vocabulary.
@@
 Tokenization is a fundamental preprocessing step in Natural Language Processing (NLP) that splits text into smaller units called tokens (e.g., words or characters), which form the input to downstream tasks.
-Over the years, various tokenization techniques have been developed (Sennrich, Haddow, and Birch 2016; Kudo and Richardson 2018; Devlin et al.
-2019).
-Among these, Byte-Pair Encoding (BPE) (Sennrich, Haddow, and Birch 2016) has become one of the most widely used.
+Over the years, various tokenization techniques have been developed (sennrich_neural_2016; kudo2018sentencepiece; devlin_bert_2019).
+Among these, Byte-Pair Encoding (BPE) (sennrich_neural_2016) has become one of the most widely used.
 Originally developed for data compression, BPE was adapted to tokenize text by iteratively merging the most frequent adjacent symbol pairs starting from individual characters until a target vocabulary size is reached.
@@
 BPE’s widespread adoption across transformer architectures has established it as a core component of LLM infrastructure.
-State-of-the-art LLMs such as GPT-4 (Achiam et al.
-2023), LLaMA (Touvron et al.
-2023a, b), Gemini (Team et al.
-2024), Claude (Anthropic 2024, 2025), and DeepSeek (Guo et al.
-2025) rely on BPE-based tokenizers to balance vocabulary efficiency and expressiveness.
+State-of-the-art LLMs such as GPT-4 (achiam2023gpt), LLaMA (touvron2023llama; touvron2023llama-1), Gemini (team2024gemini), Claude (anthropic_claude_2024; anthropic_introducing_2025), and DeepSeek (guo2025deepseek) rely on BPE-based tokenizers to balance vocabulary efficiency and expressiveness.
 However, the resulting vocabulary sizes are often large, leading to large embedding matrices and LM heads, which increase computational and memory overhead during inference.
 This scalability bottleneck has motivated research into vocabulary reduction and adaptive tokenization strategies, especially for resource-constrained deployments.
-Importantly, pruning a BPE-based tokenizer can interfere with its learned merge operations governing subword segmentation, resulting suboptimal tokenizations.
-This disruption reduces the granularity of lexical representations, ultimately impairing model performance on downstream tasks.
 2.3 Vocabulary Pruning
@@
 More recently, with the rise of small language models (SLMs), efficient vocabulary selection has once again become a pressing concern.
-Vocabulary-trimming (VT) (Ushio, Zhou, and Camacho-Collados 2023) and TextPruner (Yang, Cui, and Chen 2022) were proposed to improve efficiency and reduce model size by removing tokens irrelevant to the target language or rarely seen in downstream tasks.
+Vocabulary-trimming (ushio_efficient_2023) and TextPruner (yang_textpruner_2022) were proposed to improve efficiency and reduce model size by removing tokens irrelevant to the target language or rarely seen in downstream tasks.
 Both methods follow a static pruning strategy: they identify language-specific or task-relevant tokens from a curated corpus and prune the vocabulary accordingly.
@@
 3.1 VocabTailor Framework
-As illustrated in Fig.
-1, VocabTailor decouples vocabulary management across vocabulary-related components.
+VocabTailor (shown in Figure 1) decouples vocabulary management across vocabulary-related components.
 Unlike static vocabulary pruning that uniformly reduces the tokenizer, embedding layer, and LM head, VocabTailor treats each component based on its unique computational and storage properties.
@@
 For the embedding layer and LM head, despite their similar roles(tokenize and detokenize), a fundamental computation asymmetry exists between them.
-For the embedding layer, which relies on simple lookup operations, is naturally CPU-friendly and incurs negligible runtime overhead when offloaded to CPU memory from high-cost accelerator memory (Yu et al.
-2025).
+The embedding layer, which relies on simple lookup operations, is naturally CPU-friendly and incurs negligible runtime overhead when offloaded to CPU memory from high-cost accelerator memory yu_scaling_2025.
 Thus, we retain it fully and offload it to CPU.
-In contrast, for the LM head, which performs compute-intensive matrix multiplications, must remain on GPU for efficient inference.
+In resource-constrained deployment settings such as edge devices or platforms with unified CPU–GPU memory, even CPU-resident embeddings may impose non-trivial memory overhead.
+To address this, VocabTailor supports disk-backed embedding offloading using the Lightning Memory-Mapped Database (LMDB).
+Embedding vectors are stored as key–value entries and retrieved on demand during inference, preserving full-vocabulary access while further reducing memory footprint.
+Implementation details and analysis are in Appendix B.
+The LM head, which performs compute-intensive matrix multiplications, must remain on GPU for efficient inference.
 To optimize its memory footprint, we introduce a hybrid static-dynamic vocabulary selection strategy.
-This approach leverages the lexical locality of tokens and significantly reduces memory usage while preserving downstream task performance and enabling flexible, task-specific adaptation without model duplication.
+This approach leverages the lexical locality and significantly reduces memory usage while preserving downstream task performance and enabling flexible, task-specific adaptation without model duplication.
 This decoupled architecture addresses the key limitations of existing approaches while providing a theoretically grounded and practically efficient solution for large-scale vocabulary management.
 The detailed design and implementation of vocabulary selection strategy are presented in Sections 3.23.4.
-Figure 2: Left: Input-output lexical overlap ratio. Right: Example of input-output lexical overlap in summarization task.
 3.2 Motivation: Analysis of Lexical Locality
-To investigate the essence of efficient vocabulary selection, we analyze input-output pairs across diverse datasets aligned with various downstream tasks.
+To investigate the essence of efficient vocabulary selection, we analyze input-output pairs across diverse datasets aligned with downstream tasks.
 Our empirical analysis reveals two fundamental properties of lexical locality that motivate VocabTailor’s dynamic vocabulary selection strategy.
@@
 Input-Driven Locality: Common downstream tasks exhibit strong input-output lexical overlap—each output contains a small subset of input tokens.
-In various downstream tasks, the model output reuses tokens from the input, as shown in Fig.
-2.
+Figure 2: Left: Input-output lexical overlap ratio. Right: Example of lexical overlap in a summarization task.
+In various downstream tasks, the model output reuses tokens from the input (shown in Figure 2).
 This phenomenon is particularly pronounced in text extraction tasks (e.g., span-based QA or named entity recognition), where output tokens are typically a subset of the input tokens.
@@
 Each input introduces unique tokens, making the aggregate input vocabulary much larger than individual requirements.
-Crucially, the remaining output tokens—those not found in the corresponding input—form a relatively small, task-specific set $\mathcal{T}$ that captures the essential generation patterns for the task.
+The remaining output tokens—those not found in the corresponding input—form a relatively small, task-specific set $\mathcal{T}$ that captures the essential generation patterns for the task.
 Formally, let $\mathcal{I}$ denote the set of all input tokens, $\mathcal{I}_{i}$ the set of input tokens for instance $i$ in a dataset, $\mathcal{O}$ the set of all output tokens, and $\mathcal{T}\subset\mathcal{O}$ the set of all task-specific tokens in outputs.
-We observe that:
-$\mathcal{I}=\bigcup_{i}\mathcal{I}_{i}\gg\mathcal{I}_{i}\;\text{and }\mathcal{O}>\mathcal{T}$
-These observations expose the fundamental inefficiency of static vocabulary pruning that operates on the union $\mathcal{I}\cup\mathcal{O}$ for every example.
+We observe:
+$|\mathcal{I}|=\left|\bigcup_{i}\mathcal{I}_{i}\right|\gg|\mathcal{I}_{i}|\;\text{and }|\mathcal{O}|>|\mathcal{T}|$
+These observations expose the fundamental inefficiency of static vocabulary pruning hat operates on the union $\mathcal{I}\cup\mathcal{O}$ for every example.
 VocabTailor exploits this lexical locality by dynamically constructing active per-example vocabularies using only per-example input tokens $\mathcal{I}_{i}$ and a compact task-specific set $\mathcal{T}$:
-$\mathcal{I}\bigcup\mathcal{O}\gg\mathcal{I}_{i}\bigcup\mathcal{T}$
+$\left|\mathcal{I}\bigcup\mathcal{O}\right|\gg\left|\mathcal{I}_{i}\bigcup\mathcal{T}\right|$
 This dynamic targeting of the much smaller $\mathcal{I}_{i}\cup\mathcal{T}$ enables substantial memory savings without compromising generation quality, as it only retains the tokens necessary for each inference instance.
@@
 VocabTailor utilizes lexical locality through a hybrid architecture that combines dynamic runtime adaptation with static offline optimization, enabling efficient vocabulary management without sacrificing generation flexibility.
-Dynamic Selection (Runtime Behavior)
-At the start of each inference instance, VocabTailor identifies the unique input token set $\mathcal{I}_{i}$ from the input text and selectively loads the corresponding LM head weight vectors from CPU main memory to GPU.
+3.3.1 Dynamic Selection (Runtime Behavior)
+At the start of each inference instance, VocabTailor identifies the unique input token set $\mathcal{I}_{i}$ from the input text and selectively loads the corresponding LM head weight vectors from CPU to GPU.
 This selective loading exploits the input-driven locality principle, ensuring that only input-relevant vocabulary components are active during generation.
-Crucially, we overlap this memory transfer with the model’s forward propagation, leveraging GPU-CPU parallelism to mitigate runtime overhead.
-Static Selection (Offline Construction)
+Beyond selecting input-related tokens at runtime, VocabTailor must efficiently extend the LM head on the GPU.
+A naive implementation that repeatedly concatenates static and dynamic weights and instantiates new linear layers incurs unnecessary CPU–GPU transfers and module creation overhead.
+To mitigate this, VocabTailor utilizes the characteristics of input-related tokens and implements a pre-allocated buffer to amortize LM head construction across inputs.
+Implementations and performance evaluation are detailed in Appendix A.
+3.3.2 Static Selection (Offline Construction)
 The static component maintains a compact, task-specific core vocabulary $\mathcal{T}$ that captures essential output tokens independent of input context.
 While simple frequency-based filtering proves inadequate due to noise from typos and multilingual interference in broad corpora, we introduce a theoretically grounded filtering pipeline (Algorithm 1) that constructs $\mathcal{T}$ through fine-grained analysis.
+3.4 Fine-grained Construction of the Static Task Vocabulary
+Our static vocabulary construction algorithm (Algorithm 1) addresses the challenge of isolating task-essential tokens via a three-stage filtering process:
+Input-Aware Filtering.
+We first exclude all input tokens from the candidate vocabulary, isolating tokens the model must generate without input cues (e.g., function keywords in code generation, discourse markers in summarization).
+This step directly implements task-driven locality by identifying the irreducible core $\mathcal{T}$ that cannot be derived from the input context.
+Language-Specific Filtering.
+To suppress noise in mixed-language scenarios (e.g., code datasets with multilingual comments), we apply Unicode block analysis to retain only tokens belonging to the target language family.
+This heuristic-based approach effectively handles cross-lingual interference while preserving task-relevant vocabulary.
+Tolerance Filtering.
+When additional vocabulary reduction is required, we formulate vocabulary selection similar to a pruning problem.
+We define the impact metric $df(v)$ as the document frequency—the fraction of instances where token $v$ appears in the ground truth.
+Tokens are sorted by ascending $df(v)$ and iteratively removed until the cumulative document frequency of pruned tokens reaches a user-defined tolerance threshold $\tau$.
+This threshold bounds the worst-case performance drop, where $\tau=0.01$ ensures that at most 1% of profiling samples lose a critical token.
+VocabTailor’s hybrid approach bridges theoretical rigor (exploiting lexical locality principles) with practical efficiency (optimizing GPU-CPU memory hierarchy).
+The static-dynamic decomposition enables substantial memory savings while preserving the model’s generational capabilities, providing a scalable solution for deploying SLMs across diverse applications with controllable efficiency-accuracy trade-offs.
+4 Experiments
+4.1 Settings
+4.1.1 Tasks and datasets
+[…]
+4.1.2 Evaluation metrics
+We include sacreBLEU (post_call_2018), METEOR (banerjee_meteor_2005), and COMET (rei_comet_2020) for machine translation.
+For summarization, we use Rouge-1, Rouge-2, and Rouge-L scores (lin_rouge_2004).
+Pass@1 (chen2021evaluating) is used for code completion.
+We use F1 score for information extraction, and accuracy for math problem solving.
+These metrics are standard in the field and provide robust measures of model performance across the target tasks.
+For efficiency evaluation, we report Time to First Token (TTFT), Time Per Output Token (TPOT), Token Per Second (TPS), and Peak VRAM usage.
+4.1.3 Models
+For machine translation, we use Qwen3-1.7B (yang2025qwen3).
+We employ Llama 3.2 3B for summarization and Llama 3.2 1B for information extraction (dubey2024llama).
+For summarization, the base model is fine-tuned for a better base performance.
+For code completion, we choose deepseek-coder-1.3b-base (guo2024deepseek), and for math problem solving, we apply rho-math-1b-interpreter-v0.1 (lin2024rho).
+For efficiency evaluation, we use Qwen3-0.6B (yang2025qwen3).
+4.1.4 Baselines and Other Settings
+We compare our method (VocabTailor) with the original model (Original) and static vocabulary pruning (VP).
+Static vocabulary pruning follows the common routines of corpus-based filtering.
+For a fair comparison, VP and VocabTailor use the same profiling corpra.
+We set the tolerance threshold $\tau=0.01$ on all tasks for VocabTailor.
+As a hybrid dynamic-static framework, the vocabulary size of VocabTailor model varies in each single inference.
+Here we report the average vocabulary size for each downstream task.
+For all models, we set the temperature to 0 to avoid randomness.
+For efficiency evaluation, we use the Qwen3-0.6B model along with 100 prompts from the machine translation (English-to-Chinese) task.
+We report the latency and VRAM usage on multiple devices, including NVIDIA A100 GPU, Apple Silicon M1 Pro, and Jetson Orin Nano Super.
+4.2 Results
+Task Model Vocabulary (% Original) Metric
+Machine Translation Original 151,643 13.48/12.00/81.16
+[…]
+VocabTailor (Ours) 18,874 + [40] (12.47%) 15.39/12.69/81.44
+Machine Translation Original 151,643 24.33/48.95/73.87
+[…]
+VocabTailor (Ours) 24,185 + [14] (15.96%) 21.13/46.97/75.49
+Summarization Original 128,000 0.36/0.15/0.29
+VP 59,613 (46.57%) 0.36/0.15/0.29
+VocabTailor (Ours) 36,332 + [15] (28.40%) 0.36/0.15/0.29
+Code Completion Original 32,000 54.10%
+VP 24,888 (77.78%) 8.12%
+VocabTailor (Ours) 3,521 + [58] (11.18%) 53.87%
+Information Extraction Original 128,000 38.40
+VP 49,106 (38.36%) 2.58
+VocabTailor (Ours) [105] (0.08%) 62.73
+Math Problem Solving Original 32,000 88.40
+VP 10,300 (32.19%) 87.80
+VocabTailor (Ours) 5,135 + [14] (16.09%) 88.40
+Table 1: Results on machine translation (sacreBLEU/METEOR/COMET), summarization (Rouge-1/Rouge-2/Rouge-L), code completion (Pass@1), information extraction (F-1), and math problem solving (Accuracy), including the vocabulary size and the ratio to the original model (%).
+For VocabTailor, vocabulary size consists of task-specific tokens and the average number of dynamic tokens highlighted in brackets.
+The best results are in bold characters.
+As shown in Table 1, VocabTailor consistently achieves substantial vocabulary reduction while maintaining or even improving task performance compared with the original model, outperforming static pruning in nearly all cases.
+4.2.1 Qualitative Evaluation
+Machine Translation.
+In English-to-Chinese translation, VocabTailor achieves the best results (SacreBLEU 15.39, METEOR 12.69, and COMET 81.44), while using only 12% of the full vocabulary.
+These improvements are notable given the inherent difficulty of vocabulary pruning in high-character-set languages like Chinese.
+The results suggest that VocabTailor effectively retains the tokens most essential for both surface-level fluency and deeper semantic adequacy.
+In English-to-Italian translation, VocabTailor attains the highest COMET (75.49), indicating strong preservation of semantic and contextual meaning.
+It retains only 16% of the original vocabulary, yet performs competitively on SacreBLEU (21.13) and METEOR (46.97), with modest drops compared to the unpruned model.
+In contrast, VP removes less vocabulary (43%) but still causes moderate degradation across all metrics, suggesting that static pruning may eliminate rare but task-relevant tokens.
+VocabTailor’s ability to adapt to Italian’s rich inflectional morphology further reinforces the general applicability of the method.
+Summarization.
+For summarization, all three models yield identical ROUGE scores, indicating no difference in summary quality.
+However, VocabTailor achieves this with just 28% of the original vocabulary, 18% fewer tokens than VP.
+These results demonstrate that aggressive vocabulary reduction is possible without compromising output quality when applying a hybrid dynamic-static vocabulary selection.
+Code Completion.
+On the SAFIM benchmark with deepseek-coder-1.3b-base, VP causes a dramatic performance drop: Pass@1 falls from 54.10% to 8.12%, despite a modest vocabulary reduction to 78% of the original size.
+This highlights a key limitation of corpus-based pruning for code: essential elements such as variable names and identifiers may be discarded if they appear infrequently in the profiling corpus.
+In contrast, VocabTailor retains only 11% of the original size and achieves a high Pass@1 of 53.87%.
+These findings underscore the robustness of our input- and task-specific vocabulary selection strategy for generation-intensive tasks such as code synthesis.
+Model Device Avg. TTFT Avg. TPOT TPS Peak VRAM
+(ms) (ms) (token/s) (GB)
+Original Apple Silicon M1 Pro 23.58 75.10 18.82 1.11
+Jetson Orin Nano Super 4.40 130.80 7.68 1.17
+NVIDIA A100 1.74 24.21 41.86 1.17
+VocabTailor Apple Silicon M1 Pro 25.73 62.03 20.27 0.85
+Jetson Orin Nano Super 20.62 127.22 7.93 0.91
+NVIDIA A100 3.71 24.30 41.65 0.91
+Table 2: Performance comparison between the Original model and VocabTailor across different devices.
+Information Extraction.
+For information extraction tasks, we evaluate on the SQuAD dataset using Llama 3.2 1B.
+VocabTailor achieves the most striking result: with just 0.08% of the original vocabulary retained (and no static tokens at all), it attains an F1 score of 62.73, outperforming both the Original (38.40) and VP (2.58) by large margins.
+This result stems from the nature of extractive tasks, where the output vocabulary is typically a subset of the input.
+Because VocabTailor preserves input tokens dynamically, it retains all the necessary vocabulary for accurate extraction.
+VP’s poor performance suggests that static, corpus-based pruning is poorly suited for tasks where input-output overlap is high, whereas VocabTailor is especially effective.
+Math Word Problem Solving.
+For symbolic reasoning, we evaluate on a math problem-solving task.
+The unpruned model achieves a score of 88.40, which is fully preserved under VocabTailor, even though it reduces the vocabulary to just 16% of the Original.
+VP, while also reducing vocabulary size to 32%, causes a slight performance drop to 87.80.
+This shows that even in tasks requiring high-precision symbolic handling, our approach remains effective.
+Across five distinct tasks, VocabTailor consistently demonstrates strong performance while substantially reducing vocabulary size.
+In many cases, it matches or even exceeds the performance of the unpruned model.
+Compared to static pruning, which often compromises accuracy, VocabTailor reduces vocabulary more aggressively (up to 99%) without sacrificing model quality.
+These findings support that the dynamic vocabulary selection is a practical and efficient approach to deploying SLMs in resource-constrained environments.
+4.2.2 Efficiency Evaluation
+We evaluate the inference efficiency of VocabTailor compared to the original small language model across diverse hardware platforms, including edge devices (Apple Silicon M1 Pro and Jetson Orin Nano Super) and a high-end GPU (NVIDIA A100).
+Table 2 shows that VocabTailor consistently reduces peak VRAM usage by 22-23% (from 1.11 GB to 0.85 GB for M1 Pro chip and 1.17 GB to 0.91 GB for NVIDIA GPUs), highlighting its effectiveness in alleviating memory bottlenecks for vocabulary-related components on resource-constrained devices.
+In the decoding phase, VocabTailor achieves modest improvements in Time Per Output Token (TPOT) on lower-power hardware (up to 17% faster on M1 Pro) with negligible impact on the A100, owing to reduced computation in the dynamically trimmed LM head.
+Time to First Token (TTFT) increases on all devices, due to overhead from on-demand vocabulary loading during the prefill phase.
+However, as the runtime is dominated by decoding, VocabTailor has comparable TPS to the original model.
+4.3 Ablation Study
+We conduct an ablation study on the SAFIM dataset using deepseek-coder-1.3b-base to analyze the contribution of individual components in VocabTailor.
+We examine the roles of static and dynamic vocabulary components, as well as the effectiveness of our multi-stage filtering strategy.
+Our results show that neither static nor dynamic vocabularies alone are sufficient for optimal performance: using only dynamic (input-related) tokens leads to a substantial accuracy drop, while static-only vocabularies perform better but still underperform the hybrid static-dynamic strategy.
+This demonstrates that task-specific coverage and input-specific adaptation are complementary and jointly necessary.
+We further find that input-aware and language-specific filtering can significantly reduce the static vocabulary without degrading performance, and in some cases slightly improve accuracy by removing irrelevant tokens.
+Finally, tolerance-based filtering enables aggressive vocabulary compression with graceful degradation, offering a flexible trade-off between accuracy and memory footprint.
+Detailed experimental settings, quantitative results, and analysis are provided in Appendix E.
+5 Conclusion
+In this paper, we propose a flexible and efficient vocabulary selection framework effectively reduce memory usage during SLM inference.
+By identifying and leveraging lexical locality together with the computation asymmetry, our method reduce up to 99% in the memory usage of vocabulary-related components of SLM while maintain the performance on representative downstream tasks.
+Limitations
+The proposed framework presented in this paper is only explored on language downstream tasks of SLMs.
+This method may be extended and applied to VLMs, ALMs, and MLLMs for memory-efficient inference on image/video understanding and audio generation tasks.
+The method focuses on the memory reduction of SLMs.
+The method can be applied to LLM, but the reduction of memory will be limited.
+Appendix A Alternative Designs for Dynamic LM Head Construction
+Figure 3: Comparison of dynamic LM head construction workflows.
+Left panel (Vanilla): The naive re-allocation approach where weights are concatenated and a new Linear module is initialized per request, incurring high movement and initialization costs.
+Middle panel (SplitLinear): A decoupled architecture where static and dynamic weights form independent Linear modules ($M_{1},M_{2}$), allowing the static part to be pre-initialized on the GPU.
+Right panel (PreAlloc): Our optimized strategy using a pre-warmed GPU buffer.
+Static weights remain stationary while dynamic weights are copied into a zero-initialized buffer, eliminating module re-initialization and minimizing memory movement overhead.
+The naive (or vanilla) implementation of VocabTailor (Figure 3, left panel) dynamically constructs the LM head by selecting the input-related weights, concatenating them with the task-related static weights, onloading the weights from CPU to GPU, and finally creating a new Linear module for the concatenated LM head.
+This approach has two limitations.
+First, to avoid the VRAM peak, each time the concatenation needs to be done on the CPU.
+This results in the frequent CPU-GPU movement of the task-related static part of the LM head weights, which is highly costly for tasks requiring multi-round interactions or multiple calls to SLMs.
+Second, we need to initialize a new Linear module each time the dynamic LM head is created, and the creation of a linear module introduces extra latency overhead during the prefill stage.
+To alleviate these latencies and overheads, we propose two alternative approaches: SplitLinear and PreAlloc.
+A.1 SplitLinear
+In the SplitLinear design (Figure 3, middle panel), the static and dynamic parts of the LM head form Linear modules ($M1$ and $M2$) separately.
+Since the creation of the linear module of the static part can be done in the initialization process, only the creation of the dynamic part needs to be processed during the prefill stage.
+This design avoids weight concatenation, and the static part of the LM Head can be initialized on the GPU before inference.
+During the inference process, when the LM head is called, the input is passed through $M1$ and $M2$ independently, and their respective outputs are concatenated to form the final logit.
+Due to the nature of matrix multiplication, the forward of $M1$ and $M2$ does not interfere with each other, but two GEMMs are required per forward pass.
+A.2 PreAlloc
+To minimize the frequency of linear module creation, we investigate the number of unique input-related tokens across different downstream tasks.
+We observe that the number of unique input-related tokens is relatively small, and the distribution is concentrated.
+Among all five downstream tasks (machine translation, summarization, code completion, information extraction, and math problem solving), the average number of new unique input-related tokens per request remains below 128.
+Based on this insight, we pre-allocate a small memory space for the dynamic LM head weight.
+As shown in the right panel of Figure 3, we created a Linear module on the GPU in advance, where the weight tensor consists of two parts.
+The first part is the static weights related to the task, which are loaded before inference.
+The rest of the weight tensor is a zero-initialized buffer.
+For each input, the dynamic weights related to the input are copied into the buffer.
+As long as the dynamic weights fit within the buffer size, no concatenation and no new linear module creation are required.
+In this process, the static weights remain on the GPU without movement, so the memory movement overhead is minimized.
+In the infrequent event that the size of dynamic weights exceeds the current buffer size, a new Linear module with a larger buffer size is created in a way similar to VocabTailor’s naive implementation, thereby significantly reducing the expected latency overhead.
+A.3 Comparison of Dynamic LM Head Construction Strategies
+Both strategies are included as inference options in VocabTailor because they offer different advantages and disadvantages tailored to various downstream tasks.
+The SplitLinear approach is well-suited for downstream tasks with diverse input-related vocabularies but relatively short outputs, such as the information retrieval task, because it provides instant updates and consistent latency regardless of the number of dynamic input-related tokens.
+However, multiple GEMMs are performed in the forward pass.
+Conversely, the PreAlloc approach is ideal for downstream tasks with relatively small and stable input-related vocabularies, such as code completion or math problem solving, where the output is primarily composed of task-related terms and only a small portion of dynamic input-related vocabularies like variable names.
+In such scenarios, pre-allocation effectively eliminates both memory movement and module initialization overhead.
+A.3.1 Experimental Setup
+We test our proposed methods in the machine translation (English-to-Chinese) task using a subset of 100 examples from the WMT24++ dataset deutsch2025wmt24++.
+We use the Qwen3-0.6B yang2025qwen3 as the base model.
+All experiments are conducted on a single NVIDIA A100 GPU.
+We compared four configurations: (1) Original, the original model with standard full-vocabulary; (2) VT (Vanilla), the VocabTailor naive implementation with dynamic CPU-GPU weight concatenation; (3) VT (SplitLinear), which utilizes dual GEMM operations to bypass concatenation; and (4) VT (PreAlloc), which employs a pre-warmed GPU buffer of size 128.
+Performance is measured across 100 prompts to capture statistical distributions of latency metrics including Time to First Token (TTFT), Time Per Output Token (TPOT), and end-to-end (E2E) latency.
+We also report throughput metrics including Token Per Second (TPS), Request Per Second (RPS), and average output length.
+Memory usage is monitored via the PyTorch CUDA memory management interface to report peak VRAM consumption.
+A.3.2 Results
+Tables 3 and 4 compare the original model against three VocabTailor (VT) variants, revealing clear trade-offs between initialization overhead, steady-state decoding efficiency, and memory utilization.
+Model TTFT (ms) TPOT (ms) E2E Latency (s)
+Mean P50 P90 P99 Mean P50 P90 P99 Mean P50 P90 P99
+Original 1.74 0.79 0.90 2.07 24.21 23.64 24.09 25.17 1.57 1.56 2.44 4.12
+VT (Vanilla) 124.29 123.22 124.63 137.43 23.99 23.47 23.60 24.74 1.66 1.66 2.47 3.91
+VT (SplitLinear) 123.86 123.38 124.55 129.62 24.52 24.05 24.19 28.87 1.69 1.69 2.53 3.98
+VT (PreAlloc) 3.71 2.63 3.35 17.20 24.30 23.82 24.32 25.96 1.56 1.57 2.38 3.91
+Table 3: Latency distribution for different dynamic LM head construction approaches on a machine translation (English-to-Chinese) task ($N=100$).
+We report Time to First Token (TTFT), Time Per Output Token (TPOT), and end-to-end (E2E) latency.
+Percentiles (Mean, P50, P90, P99) are included to characterize tail latency and model stability under varying prompt complexity.
+Model TPS RPS Avg. Output Peak VRAM
+(token/s) (req/s) Length (GB)
+Original 41.86 0.64 65.73 1.17
+VT (Vanilla) 42.31 0.56 64.84 0.91
+VT (SplitLinear) 41.34 0.55 64.84 0.91
+VT (PreAlloc) 41.65 0.55 64.84 0.91
+Table 4: Throughput and resource utilization for different dynamic LM head construction approaches on a machine translation (English-to-Chinese) task ($N=100$).
+Decode throughput is measured in Tokens Per Second (TPS), computed as total generated tokens divided by total decoding time.
+Request throughput (RPS), average output length, and peak GPU memory consumption are also reported.
+Latency behavior.
+The Original achieves the lowest Time to First Token (TTFT), with a mean of 1.74 ms, reflecting its fully static vocabulary and absence of runtime adaptation.
+In contrast, the Vanilla VT incurs a substantial TTFT increase (mean 124.29 ms), which can be attributed to dynamic vocabulary construction and associated runtime bookkeeping.
+Notably, this overhead is largely confined to the prefill phase: Time Per Output Token (TPOT) remains comparable across all models ($\approx$24 ms), indicating that VT does not degrade steady-state decoding once generation begins.
+Among different dynamic head construction approaches, PreAlloc is particularly effective at mitigating TTFT overhead, reducing mean TTFT to 3.71 ms, within the same order of magnitude as Original, while preserving the benefits of VT.
+Tail latency analysis further supports this observation: PreAlloc substantially improves P99 TTFT (17.20 ms vs.
+137.43 ms for Vanilla VT), suggesting improved stability under varying prompt complexities.
+In contrast, SplitLinear does not materially reduce TTFT, as it still requires the dynamic creation of the module $M2$.
+End-to-end latency.
+Despite large TTFT differences, end-to-end (E2E) latency remains broadly similar across models.
+All VT variants exhibit mean E2E latency within 1.56–1.69 s, comparable to the Original (1.57 s).
+This indicates that, for typical MT tasks with non-trivial output lengths, TTFT overhead is amortized over decoding, and overall user-perceived latency is dominated by generation rather than initialization.
+Latency decomposition.
+Figure 4 illustrates the breakdown of total E2E latency consumed by the prefill versus decoding stages.
+In the Vanilla and SplitLinear implementations, the prefill stage—which is effectively instantaneous in the baseline ($0.11\%$)—surges to over $7.3\%$ of the total execution time.
+By utilizing a pre-warmed GPU buffer, the pre-allocation strategy successfully reduce the prefilling time, returning the prefill stage to just $0.24\%$ of total latency.
+Figure 4: Latency decomposition across different dynamic LM head construction approaches.
+PreAlloc (bottom) successfully reduces the total prefill time from 12.43s (Vanilla) back to 0.37s, effectively matching the latency profile of the original model.
+Throughput and resource efficiency.
+As shown in Table 4, decode throughput (TPS) remains effectively consistent across all models (41–42 tokens/s), confirming that VT and its extensions do not introduce steady-state performance regressions.
+Request-level throughput, measured by Request Per Second (RPS), is slightly lower for VT-based models, which aligns with their increased per-request initialization cost.
+In terms of memory footprint, VT variants consistently reduce peak VRAM usage (0.91 GB vs.
+1.17 GB for Original), validating the effectiveness of vocabulary decoupling and offloading.
+Overall trade-offs.
+Taken together, these results highlight that VocabTailor introduces a clear TTFT–memory trade-off: substantial VRAM savings with minimal impact on decode throughput, at the cost of higher initialization latency.
+Among the different head construction strategies, PreAlloc offers the most favorable balance, largely eliminating TTFT penalties while preserving VRAM savings and decode efficiency.
+SplitLinear  while maintaining comparable throughput and memory characteristics, provides limited benefit for latency-sensitive scenarios.
+Appendix B Offload Embedding Lookup to LMDB
+Figure 5: Overview of the VocabTailor framework with disk-backed embedding offloading.
+The embedding layer is replaced with a custom embedding layer that retrieves the corresponding tokens from the LMDB when the forward is called.
+In VocabTailor, model embeddings are offloaded to the CPU to reduce GPU memory usage.
+However, on devices with limited CPU memory or unified CPU-GPU memory architectures, this design alone does not sufficiently mitigate memory consumption.
+Such constraints are prevalent in edge devices, where SLMs are frequently deployed, as matrix multiplications can still be executed on the CPU within acceptable latency.
+In this context, VocabTailor’s core principle—decoupling and offloading embedding weights to a lower storage hierarchy—remains effective for two primary reasons: (1) embedding lookup operations are computationally lightweight, and (2) the dominant inference latency arises from the forward passes, particularly tensor multiplications, making data movement a secondary bottleneck.
+Building on this principle and inspired by prior work on memory-mapped embedding storage (yu_scaling_2025), we implement disk-based embedding offloading using the Lightning Memory-Mapped Database (LMDB).
+The embedding weights are serialized as key-value pairs, with token indices as keys and the corresponding embedding vectors as values.
+A disk-backed LMDB database is constructed to store these pairs, enabling on-demand retrieval of embeddings during inference.
+The extended framework is illustrated in Figure 5.
+B.1 Experimental Setup
+We evaluate the effect of embedding offloading for the machine translation (English-to-Chinese) task on a subset of 100 examples from the WMT24++ dataset deutsch2025wmt24++.
+We use the Qwen3-0.6B yang2025qwen3 as the base model.
+All experiments are conducted on a single NVIDIA A100 GPU.
+Building on our previous results identifying PreAlloc as the most efficient dynamic LM head construction strategy, we compare three configurations that progressively incorporate embedding offloading: (1) Original, the original model with standard full-vocabulary; (2) VT (PreAlloc), which offloads embedding weights to CPU and employs a pre-warmed GPU buffer of size 128; and (3) VT (PreAlloc) + DiskEmb, which further offloads embeddings to disk-backed LMDB storage.
+Performance is measured across 100 prompts to capture statistical distributions of latency metrics including Time to First Token (TTFT), Time Per Output Token (TPOT), and end-to-end (E2E) latency.
+We also report throughput metrics including Token Per Second (TPS), Request Per Second (RPS), and average output length.
+Memory usage is monitored via the PyTorch CUDA memory management interface to report peak VRAM consumption.
+B.2 Results
+Table 5 presents a comparison between the three configurations.
+The latter two configurations share the same dynamic LM head construction strategy and differ only in how embedding weights are stored and accessed.
+Model TTFT (ms) TPOT (ms) E2E Latency (s)
+Mean P50 P90 P99 Mean P50 P90 P99 Mean P50 P90 P99
+Original 1.74 0.79 0.90 2.07 24.21 23.64 24.09 25.17 1.57 1.56 2.44 4.12
+VT (PreAlloc) 3.71 2.63 3.35 17.20 24.30 23.82 24.32 25.96 1.56 1.57 2.38 3.91
+VT (PreAlloc) + DiskEmb 4.58 3.64 4.10 12.23 24.46 24.01 24.27 25.50 1.57 1.58 2.39 3.82
+Table 5: Latency distribution for three inference configurations on a machine translation (English-to-Chinese) task ($N=100$), comparing the effect of embedding offloading.
+We report Time to First Token (TTFT), Time Per Output Token (TPOT), and end-to-end (E2E) latency.
+Percentiles (Mean, P50, P90, P99) are included to characterize tail latency and model stability under varying prompt complexity.
+Model TPS RPS Avg. Output Weights on CPU Peak VRAM
+(token/s) (req/s) Length (GB) (GB)
+Original 41.86 0.64 65.73 0 1.17
+VT (PreAlloc) 41.65 0.55 64.84 0.28 0.91
+VT (PreAlloc) + DiskEmb 41.44 0.51 64.78 0 0.91
+Table 6: Throughput and resource utilization for three inference configurations on a machine translation (English-to-Chinese) task ($N=100$), comparing the effect of embedding offloading.
+Decode throughput is measured in Tokens Per Second (TPS), computed as total generated tokens divided by total decoding time.
+Request throughput (RPS), average output length, weights on CPU, and peak GPU memory consumption are also reported.
+Latency behavior.
+Compared to the Original, VT (PreAlloc) increases mean TTFT from 1.74 ms to 3.71 ms, reflecting the additional overhead of dynamic LM head construction and CPU-based embedding lookups.
+Introducing disk-based offloading further increases TTFT slightly to 4.58 ms, as embedding vectors are retrieved from disk rather than CPU memory.
+Importantly, this overhead remains small in absolute terms.
+Tail latency analysis reveals a different trend.
+While VT (PreAlloc) exhibits a higher P99 TTFT (17.20 ms), DiskEmb configuration reduces P99 TTFT to 12.23 ms, indicating more stable prefill latency.
+This suggests that disk access introduces predictable overhead without amplifying variance across inputs.
+TPOT remains comparable across all configurations, with mean TPOT values clustered around 24 ms.
+This confirms that embedding offloading—whether to CPU or LMDB—does not affect steady-state decoding, which is dominated by transformer forward passes.
+End-to-end (E2E) latency remains broadly consistent across configurations.
+Despite differences in TTFT, both VT-based models achieve comparable mean and slightly lower tail E2E latency relative to the Original, indicating that prefill overhead is amortized over the generation process for typical output lengths.
+Latency decomposition.
+Figure 6 illustrates the breakdown of total end-to-end (E2E) latency consumed by the prefill stage versus the decode stage.
+In the vanilla configuration, the prefill stage is nearly instantaneous, accounting for only 0.11% of total execution time.
+Transitioning to VT (PreAlloc) increases this proportion slightly to 0.24%.
+When moving further to DiskEmb, the prefill percentage still remains remarkably low at 0.30%.
+This demonstrates that while retrieving embedding vectors from disk-based storage (LMDB) is technically more time-consuming than CPU memory access, the impact on the overall latency profile is negligible.
+In all cases, the decoding stage continues to dominate over 99.7% of the runtime.
+Figure 6: Latency decomposition of three inference configurations.
+We report the total end-to-end (E2E) latency consumed by the prefill (initialization) stage versus the decoding (generation) stage.
+While offloading embeddings to the CPU or LMDB increases the prefill proportion relative to the Original ($0.11\%$), the impact remains marginal ($0.24\%$ for VT PreAlloc and $0.30\%$ for DiskEmb), ensuring that the latency remains dominated by decoding.
+Throughput and resource efficiency.
+Table 6 reports decode throughput and resource usage.
+Decode TPS remains stable across configurations (around 42 tokens/s).
+This supports the observation that offloading embedding does not degrade decoding performance.
+RPS is slightly lower for VT variants, consistent with their higher per-request initialization overhead.
+Peak GPU memory is reduced from 1.17 GB (Original) to 0.91 GB for both VT-based configurations, reflecting the benefit of pre-allocating LM head weights on GPU and offloading embeddings.
+VT (PreAlloc) has embedding weights on the CPU, which consumes 0.28 GB of memory.
+With DiskEmb, the embedding is evicted to disk, eliminating the CPU memory overhead.
+Summary.
+Overall, the results confirm that embedding offloading is compatible with VocabTailor and provides additional memory savings.
+Both CPU-based and disk-backed offloading improve GPU memory efficiency relative to the original model without compromising decoding performance, making VocabTailor suitable for memory-constrained or resource-limited environments.
+Appendix C Support Both Tied and Non-Tied Embedding Architectures
+In large language models, the embedding layer and the LM head share the same weight dimension.
+The hidden states after going through the LM head and the input tokens before feeding into the transformer layers are considered in a similar representation space.
+Thus, to reduce the total number of parameters, some models would share the weight of the embedding and the LM head.
+Weight tying effectively reduces model size and inference-time memory consumption and is therefore widely adopted in small language models (SLMs).
+However, a portion of SLMs still retains non-tied embeddings to keep higher expressiveness.
+Figure 7: Comparison of VocabTailor workflows.
+Left: In tied architectures, LM head weights are reused for decoding to eliminate memory movement.
+ID mapping is deferred until the final tokenizer stage.
+Right: In non-tied architectures, only the LM head is reduced, so an explicit ID mapping is required after sampling to align the reduced LM head and the full CPU embedding layer.
+VocabTailor supports both tied and non-tied embedding architectures, with their respective inference workflows illustrated in Figure 7.
+For tied embedding models, token lookup during the prefill stage is performed using the embedding weights stored on the CPU.
+During decoding, because the embedding and LM head weights are shared, we can directly use the reduced LM head weights to replace the embedding weights without extra memory usage, thereby eliminating memory movement overhead.
+Moreover, since the LM head is already reduced, embedding lookup can be performed without an ID mapping between the original token index and the reduced version token index.
+Such mapping is deferred until generation completes, before tokenizer decoding.
+For the non-tied embedding models, the embedding layer remains on the CPU and is used for both prefill and decode stages.
+Since only the LM head is reduced, the token indices produced by the reduced LM head after the softmax and sampling operation must be mapped back to the original indices before calling the embedding layer forward to get the corresponding token vectors.
+After generation, the produced token IDs are already the same as the ones in the full vocabulary, so no additional ID mapping is required for the tokenizer decoding.
+Appendix D Profiling Strategy Algorithm
+The detailed profiling strategy is presented in Algorithm 1.
 Algorithm 1 Profiling Strategy
@@
 17:return $\mathcal{T}\leftarrow\{v\in F[-\texttt{index}:]\}$
-3.4 Fine-grained Construction of the Static Task Vocabulary
-Our static vocabulary construction algorithm addresses the challenge of isolating truly task-essential tokens through a three-stage filtering process:
-Input-Aware Filtering
-We first exclude all input tokens from the candidate vocabulary, isolating tokens the model must generate without input cues (e.g., function keywords in code generation, discourse markers in summarization).
-This step directly implements task-driven locality by identifying the irreducible core $\mathcal{T}$ that cannot be derived from the input context.
-Language-Specific Filtering
-To suppress noise in mixed-language scenarios (e.g., code datasets with multilingual comments), we apply Unicode block analysis to retain only tokens belonging to the target language family.
-This heuristic-based approach effectively handles cross-lingual interference while preserving task-relevant vocabulary.
-Tolerance Filtering
-When additional vocabulary reduction is required, we formulate vocabulary selection similar to a pruning problem.
-We define the impact metric $df(v)$ as the document frequency—the fraction of instances where token $v$ appears in the ground truth.
-Tokens are sorted by ascending $df(v)$ and iteratively removed until the cumulative document frequency of pruned tokens reaches a user-defined tolerance threshold $\tau$.
-This threshold bounds the worst-case performance drop, where $\tau=0.01$ ensures that at most 1% of profiling samples lose a critical token.
-VocabTailor’s hybrid approach bridges theoretical rigor (exploiting lexical locality principles) with practical efficiency (optimizing GPU-CPU memory hierarchy).
-The static-dynamic decomposition enables substantial memory savings while preserving the model’s generational capabilities, providing a scalable solution for deploying SLMs across diverse applications with controllable efficiency-accuracy trade-offs.
-4 Experiments
-4.1 Settings
-Tasks and datasets
-[…]
-Evaluation metrics
-We include sacreBLEU (Post 2018), METEOR (Banerjee and Lavie 2005), and COMET (Rei et al.
-2020) for machine translation.
-For summarization, we use Rouge-1, Rouge-2, and Rouge-L scores (Lin 2004).
-Pass@1 (Chen et al.
-2021) is used for code completion.
-We use F1 score for information extraction, and accuracy for math problem solving.
-These metrics are standard in the field and provide a robust measure of model performance across the target task.
-Models
-For machine translation, we use Qwen3-1.7B (Yang et al.
-2025).
-For summarization and information extraction, we employ Llama 3.2 1B (Dubey et al.
-2024).
-For summarization, the base model is fine-tuned for a better base performance.
-For code completion, we choose deepseek-coder-1.3b-base (Guo et al.
-2024), and for math problem solving, we apply rho-math-1b-interpreter-v0.1 (Lin et al.
-2024).
-Baselines and Other Settings
-We compare our method (VocabTailor) with the original model (Original) and static vocabulary pruning (VP).
-Static vocabulary pruning follows the common routines of corpus-based filtering.
-For a fair comparison, VP and VocabTailor use the same profiling corpra.
-We set the tolerance threshold $\tau=0.01$ on all tasks for VocabTailor.
-As a hybrid dynamic-static framework, the vocabulary size of VocabTailor model varies in each single inference.
-Here we report the average vocabulary size for each downstream task.
-For all models, we set the temperature to 0 to avoid randomness.
-4.2 Results
-Task Model Vocabulary Metric
-[…]
-VP 64,117 (42.28%) 14.840/11.842/81.191
-VocabTailor (Ours) 18,874 + [40] (12.47%) 15.393/12.692/81.440
-[…]
-VP 65,518 (43.21%) 22.585/46.735/74.933
-VocabTailor (Ours) 24,185 + [14] (15.96%) 21.127/46.970/75.493
-Summarization Original 128,000 0.364/0.147/0.292
-VP 59,613 (90.99%) 0.364/0.147/0.292
-VocabTailor (Ours) 36,332 + [15] (55.48%) 0.364/0.147/0.292
-Code Completion Original 32,000 54.097%
-VP 24,888 (77.78%) 8.124%
-VocabTailor (Ours) 3,521 + [58] (11.18%) 53.865%
-Information Extraction Original 128,000 38.400
-VP 49,106 (38.36%) 2.580
-VocabTailor (Ours) [105] (0.08%) 62.730
-Math Problem Solving Original 32,000 88.40
-VP 10,300 (32.19%) 87.80
-VocabTailor (Ours) 5,135 + [14] (16.09%) 88.40
-Table 1: Results on machine translation (sacreBLEU/METEOR/COMET), summarization (Rouge-1/Rouge-2/Rouge-L), code completion (Pass@1), information extraction (F-1), and math problem solving (Accuracy), including the vocabulary size and the ratio to the original model (%).
-For VocabTailor, vocabulary size consists of task-specific tokens and the average number of dynamic tokens highlighted in brackets.
-The best results are in bold characters.
-As shown in Table 1, VocabTailor consistently achieves substantial vocabulary reduction while maintaining or even improving task performance compared with the original model, outperforming static pruning in nearly all cases.
-Machine Translation
-[…]
-[…]
-Summarization
-For summarization, all three models yield identical ROUGE scores, indicating no difference in summary quality.
-However, VocabTailor achieves this with just 55% of the original vocabulary, 35% fewer tokens than VP.
-These results demonstrate that aggressive vocabulary reduction is possible without compromising output quality when apply a hybrid dynamic-static vocabulary selection.
-Code Completion
-On the SAFIM benchmark with deepseek-coder-1.3b-base, VP leads to a dramatic drop in performance: Pass@1 falls from 54.097% to 8.124%, despite a relatively modest vocabulary reduction (down to 77.78% of the original size).
-This highlights a key limitation of corpus-based pruning for code: essential elements such as variable names and identifiers may be discarded if they appear infrequently in the profiling corpus.
-In contrast, VocabTailor retains only 11.18% of the vocabulary and yet achieves a high Pass@1 of 53.865%.
-These findings underscore the robustness of our input- and task-specific vocabulary selection strategy for generation-intensive tasks such as code synthesis.
-Information Extraction
-For information extraction tasks, we evaluate on the SQuAD dataset using LLaMA 3.2 1B.
-VocabTailor achieves the most striking result: with just 0.08% of the original vocabulary retained (and no static tokens at all), it attains an F1 score of 62.73, outperforming both the Original (38.4) and VP (2.58) by large margins.
-This result stems from the nature of extractive tasks, where the output vocabulary is typically a subset of the input.
-Because VocabTailor preserves input tokens dynamically, it retains all the necessary vocabulary for accurate extraction.
-The poor performance of VP suggests that static, corpus-based pruning is poorly suited for tasks where input-output overlap is high, whereas VocabTailor is especially effective.
-Math Word Problem Solving
-For symbolic reasoning, we evaluate on a math problem-solving task.
-The unpruned model achieves a score of 88.40, which is fully preserved under VocabTailor, even though it reduces the vocabulary to just 16% of the original.
-VP, while also reducing vocabulary size to 32%, causes a slight performance drop to 87.80.
-This shows that even in tasks requiring high-precision symbolic handling, our approach remains effective.
-Across five distinct tasks, VocabTailor consistently demonstrates strong performance while substantially reducing vocabulary size.
-In many cases, it matches or even exceeds the performance of the unpruned model.
-Compared to static pruning, which often compromises accuracy, VocabTailor reduces vocabulary more aggressively (up to 99%) without sacrificing model quality.
-These findings support that the dynamic vocabulary selection is a practical and efficient approach to deploying SLMs in resource-constrained environments.
-4.3 Ablation Study
+Appendix E Ablation Study in Details
 To understand the contributions of each component in our framework, we conduct a series of ablation experiments on the SAFIM dataset using the deepseek-coder-1.3b-base model for the code completion task.
 We primarily focus on evaluating the impact of different vocabulary configurations, including the static and dynamic components, as well as our proposed three-stage filtering process.
-Impact of Dynamic vs. Static Vocabulary Components
+E.1 Impact of Dynamic vs. Static Vocabulary Components
 We first evaluate the impact of the dynamic and static vocabulary components, both individually and in combination.
-As shown in Table 2, using only the dynamic part that contains tokens profiled from the specific input examples, results in a significant performance drop (Pass@1 of 36.064%).
+As shown in Table 7, using only the dynamic part that contains tokens profiled from the specific input examples results in a significant performance drop (Pass@1 of 36.06%).
 This demonstrates that input tokens alone lack the broader coverage needed for robust code completion.
-Using only the static part (task-specific tokens) achieves a Pass@1 of 52.3%.
-However, this still underperforms the full static-dynamic configuration, which achieves the best result at 53.865% with nearly the same vocabulary size.
+Using only the static part (task-specific tokens) achieves a Pass@1 of 52.30%.
+However, this still underperforms the full static-dynamic configuration, which achieves the best result at 53.87% with nearly the same vocabulary size.
 This indicates that while the static tokens carry most of the task-relevant capacity, including the dynamic tokens adds critical input-specific nuances, and their combination is essential for optimal performance.
-We also compare our static-only approach with VP: VP retains 78% of the original vocabulary—more than our approach—it results in a drastically lower Pass@1 of 8.124%.
+Model Vocabulary Pass@1
+Original 100% 54.10%
+VP 77.80% 8.12%
+Dynamic + Static ($\tau=0.01$) 11.18% 53.87%
+Dynamic only 0.81% 36.06%
+Static only ($\tau=0.01$) 11.00% 52.30%
+Table 7: Comparison of the dynamic and static components in VocabTailor.
+We also compare our static-only approach with VP: VP retains 78% of the original vocabulary—more than our approach—it results in a drastically lower Pass@1 of 8.12%.
 This stark contrast underscores a key insight: while both VP and our static-only setup are static, VP’s direct modification of the tokenizer and embeddings damages input representations, leading to severe degradation.
 In contrast, our method retains the full tokenizer and embedding, thereby preserving representational integrity and maintaining high performance even with significantly fewer tokens.
-Model Vocabulary Pass@1
-Original 100% 54.097%
-VP 77.80% 8.124%
-Dynamic + Static ($\tau=0.01$) 11.18% 53.865%
-Dynamic only 0.81% 36.064%
-Static only ($\tau=0.01$) 11.00% 52.300%
-Table 2: Comparison of the dynamic and static components in VocabTailor.
-Impact of Input-aware and Language-specific Filtering
+E.2 Impact of Input-aware and Language-specific Filtering
 As discussed earlier, VocabTailor calibrates the static vocabulary through three-stage filtering.
-As shown in Table 3, starting from the unfiltered static vocabulary (78% of the original), applying input-aware filtering (IA) reduces the size to 53% with virtually no performance loss (54.074% vs.
-54.062%).
-Adding language-specific filtering (LS) further reduces the vocabulary size to 46%, while performance slightly improves to 54.085%.
+In Table 8, starting from the unfiltered static vocabulary (78% of the Original), applying input-aware filtering (IA) reduces the size to 53% with virtually no performance loss (54.07% vs.
+54.06%).
+Adding language-specific filtering (LS) further reduces the vocabulary size to 46%, while performance slightly improves to 54.09%.
 This improvement likely stems from the removal of noisy or irrelevant tokens from multilingual corpora, allowing the model to focus on task-relevant representations.
@@
 Model Vocabulary Pass@1
-Original 100.00% 54.097%
-Dynamic + Unfiltered static 77.78% 54.074%
-Dynamic + IA 52.85% 54.062%
-Dynamic + IA + LS 45.61% 54.085%
+Original 100.00% 54.10%
+Dynamic + Unfiltered Static 77.78% 54.07%
+Dynamic + IA 52.85% 54.06%
+Dynamic + IA + LS 45.61% 54.09%
 Table 8: Comparison of input-aware filtering (IA) and language-specific filtering (LS) in VocabTailor.
-Impact of Tolerance Filtering
+E.3 Impact of Tolerance Filtering
 Lastly, we analyze the effect of tolerance filtering, which allows for further reduction of rarely activated tokens based on cumulative document frequency.
 In Table 9, we vary the tolerance threshold ($\tau$) to observe the trade-off between vocabulary size and accuracy.
-At $\tau=0$, we preserve all profiled tokens, achieving a Pass@1 of 54.085%.
+At $\tau=0$, we preserve all profiled tokens, achieving a Pass@1 of 54.09%.
 As tolerance increases, more tokens are filtered out, reducing the vocabulary to as low as 2.5% of the original size ($\tau=0.10$), with a gradual decline in performance.
-Importantly, even with $\tau=0.01$, the vocabulary shrinks to 11.18% with only 1.8% drop in Pass@1, suggesting that our method is robust to aggressive reduction.
+Importantly, even with $\tau=0.01$, the vocabulary shrinks to 11% with only 1.8% drop in Pass@1, suggesting that our method is robust to aggressive reduction.
 This highlights the flexibility of tolerance as a tuning knob to balance compression vs.
@@
   Model   Vocabulary   Pass@1
-$\tau=0$ 45.61% 54.085%
-$\tau=0.01$ 11.18% 53.865%
-$\tau=0.02$ 7.28% 53.714%
-$\tau=0.10$ 2.50% 52.277%
+  $\tau=0$   45.61%   54.09%
+  $\tau=0.01$   11.18%   53.87%
+  $\tau=0.02$   7.28%   53.71%
+  $\tau=0.10$   2.50%   52.28%
 Table 9: Comparison of different tolerance thresholds in VocabTailor.
-5 Conclusion
-In this paper, we propose a flexible and efficient vocabulary selection framework effectively reduce memory usage during SLM inference.
-By identifying and leveraging lexical locality together with the computation asymmetry, our method reduce up to 99% in the memory usage of vocabulary-related components of SLM while maintain the performance on representative downstream tasks.
-6 Limitations & Future Work
-The proposed framework presented in this paper is only explored on language downstream tasks of SLMs.
-This method may be extended and applied to VLMs, ALMs, and MLLMs for memory-efficient inference on image/video understanding and audio generation tasks.
-Custom kernel implementations and hardware-specific optimizations can be integrated into the framework for further inference optimization.
+Appendix F Data Processing and Evaluation Details
+For machine translation, we follow the best practice of zero-shot machine translation by zhang_prompting_2023, using a simple English template.
+English-to-Chinese translation template:
+[…]
+English-to-Italian translation template:
+[…]
+For summarization, the model is finetuned and evaluated using prompt shown below.
+Summarization prompt template:
+[…]
+For the code completion task, we use the evaluation script of SAFIM.
+For information extraction, we use the LM-Eval-Harness framework and set the task to ‘squad_completion‘.
+For math word problem solving, we use the Math-Eval-Harness framework and set the task to ‘mawps‘.
+Appendix G Fine-tuning Config for Summarization Baseline Model
+For summarization, all the methods are based on a finetuned Llama 3.2 3B model. We set learning rate=2e-5, batch size = 128. The model is fine-tuned using XSUM train split set for 1 epoch.